if the zeros of f(x) are 2 and 3, what are the zeros of f(x+1)?
Answer by edjones(8007)

x=2, x=3
(x-2)(x-3)=0
f(x)=x^2-5x+6=0
f(x+2)=(x+2)^2-5(x+2)+6=0
x^2+4x+4-5x-10+6=0
x^2-x=0
x(x-1)=0
x=0, x=1
.
Ed